OpenStudy (amistre64):

What are all the applications that can take advantage of a Larange Multiplier?

OpenStudy (amistre64):

lagrange even lol

OpenStudy (anonymous):

min/max

OpenStudy (amistre64):

min max is one; any others?

OpenStudy (amistre64):

and how would you go about finding a min with it?

OpenStudy (anonymous):

you can't find min/max, just extrema

OpenStudy (anonymous):

so when you find a extrema, compare with other point to see if it is min or max

OpenStudy (anonymous):

any other application Where two gradient vector are same direction

OpenStudy (amistre64):

I was trying to see if i could use it last night to solve a system of equations :) dint work as i expected lol

OpenStudy (amistre64):

f(x,y) = 3x-4y - 20 g(x,y) = 2x +y +3 F(x,y,L) = 3x-4y - 20 - L(2x +y +3) :)

OpenStudy (amistre64):

Fx = 3-2L = 0 L = 3/2 Fy = -4-L = 0 L = -4 FL = 2x+y+3 = 0 .....

OpenStudy (amistre64):

but since there is no spot for the gradients to be the same; i guess it just wont pan out

OpenStudy (anonymous):

As you can see from picture , there are no places where gradient vector are in same direction

OpenStudy (amistre64):

yep; that was my conclusions as well ;)

OpenStudy (anonymous):

That's interesting because if it were a min/max function with one plane a constraint, it wouldn't work

OpenStudy (amistre64):

even if there was a point of common gradience; the points of intersection required from a system of equations would not be condusive to a lagrange multi ... :)

OpenStudy (amistre64):

a lagrange is useful in finding the common point that 2 systems have in tangency to each other; where they softly touch ... finding secants and chords requires something else i believe

OpenStudy (anonymous):

you take the cross product of two gradient vector {3,4}x{2,1}

OpenStudy (amistre64):

also known as the determinant right? 34 21 3 - 8 = -5